A solution containing 9.15 g of calcium nitrate is added to a solution containing 9.15 g of potassium phosphate. What is the theoretical yield of the precipitate in grams? Report your answer rounded to the hundredths of a gram.
Answer #1

Molar mass of Ca(NO3) 2 = 40.08 + ( 2 \times 14.0067) + ( 6 \times 16.00) = 164.09 g / mol

Molar mass of Ca3 (PO4)2 = ( 3 \times 40.08 ) + ( 2 \times 30.97) + ( 8 \times 16.00) = 310.18 g/mol

Molar mass of K3PO4 = ( 3 \times 39.10 ) + 30.97 + ( 4 \times 16.00) = 212.27 g/ mol

Consider a precipitation reaction, 3 Ca(NO3) 2 (aq) + 2 K3PO4 (aq) \rightarrow Ca3 (PO4)2 (s) + 6 KNO 3 (aq)

From reaction , 3 mol Ca(NO3) 2 \equiv 2 mol K3PO4\equiv1 mol Ca3 (PO4)2

3 \times 164.09 g Ca(NO3) 2 \equiv 2 \times 212.97 g K3PO4

9.15 g  Ca(NO3) 2 \equiv 2 \times 212.97 \times 9.15 / 3 \times 164.09 g K3PO4

9.15 g  Ca(NO3) 2 \equiv 7.917 g K3PO4

i e 9.15 g calcium nitrate requires 7.917 g potassium phosphate. Hence, potassium phosphate is excess reactant because its amount (9.15 g) is excess than required amount ( 7.917 g)

Calcium nitrate is limiting reactant , hence mass of product will depend on mass of calcium nitrate.

We have relation,  3 mol Ca(NO3) 2 \equiv 1 mol Ca3 (PO4)2

3 \times 164.09 g Ca(NO3) 2 \equiv 310.18 g Ca3 (PO4)2

9.15 g  Ca(NO3) 2 \equiv  310.18 \times 9.15 / ( 3 \times 164.09) g Ca3 (PO4)2

9.15 g  Ca(NO3) 2 \equiv 5.76 g Ca3 (PO4)2

ANSWER : Theoretical yield of precipitate in grams = 5.76 g